OpenStudy (anonymous):

During a recent election, 46 of every 100 registered voters in a certain city voted. If there were 158,900 registered voters in that city, how many people voted? people voted in the recent election

OpenStudy (anonymous):

@roz_mar23

OpenStudy (anonymous):

\[\frac{ 46 }{ 100}=\frac{ ? }{ 158,900 }\]

OpenStudy (anonymous):

im confused?

OpenStudy (anonymous):

What times 100=158,900?

OpenStudy (anonymous):

100*1589

OpenStudy (anonymous):

so 46x1589=73,094 \[\frac{ 46 }{ 100 } =\frac{ 73,094 }{ 158,900 }\]

OpenStudy (anonymous):

46/100=.46 73094/158900=.46 So they both equal .46 so the answer is 73094

OpenStudy (anonymous):

ahhh gotchya! I get it:)